OpenStudy (anonymous):

applications of rational equations: A painter works on a job for 10 days and is then joined by an associate. Together they finish the job in 6 more days. the associate could have done the job alone in 30 days. How long would it take the painter to do the job alone?

OpenStudy (anonymous):

The associate can do the job alone in 30 days, meaning that in 1 day, he can complete 1/30 of the job. In the 6 days, when both the painter and the associate are working together, the associate completed 6/30= 1/5 of the job. The painter, in his first 10 days and the 6 days (total 16days) he done the rest of the job. He took 16 days to complete 4/5 of the job. Number of days he will take to complete the job alone, 16/4 *5 = 20 days.

OpenStudy (anonymous):

so, what if I used that in a numerical solution>?

OpenStudy (anonymous):

hmm what i typed is a solution?

OpenStudy (anonymous):

can you make it mor simpler like just put the process on how to solve it.. I know the question was soo easy but my head doesn't wanna solve math problems which is annoying >.<